OpenStudy (anonymous):

Which equation is not an identity? tan(–t) = –tan t tan t(csc t + cot t) = sec t + 1 sin tcot t = cos t cos2t − sin2t = –1

myininaya (myininaya):

let's look at the first equation what kinda function is tan? is it even or odd?

OpenStudy (cwrw238):

an identity is true for all values of the variable

OpenStudy (anonymous):

Well, tangent is an odd function.

myininaya (myininaya):

I was hoping Alg could tell us that

OpenStudy (anonymous):

Okay there are rules, Identities: \(\normalsize\color{black}{ \sin(~-t~)=-\sin(~t~) }\) \(\normalsize\color{black}{ \cos(~-t~)=\cos(~t~) }\) \(\normalsize\color{black}{ \tan(~-t~)=-\sin(~t~) }\) Don't confuse: \(\normalsize\color{red}{ \sin(~-t~) \ne\sin(~t~) }\) \(\normalsize\color{red}{ \cos(~-t~)\ne -\cos(~t~) }\) \(\normalsize\color{red}{ \tan(~-t~)=\tan(~t~) }\)

OpenStudy (anonymous):

the last one is supposed to be \(\normalsize\color{red}{ \tan(~-t~) \ne \tan(~t~) }\)